OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality: 7 - 9x - (x + 12) 25? A: x -3 B: x -3 C: -3 x D: A number line with a closed circle on negative 3 and shading to the right

OpenStudy (aripotta):

what's the inequality symbol here?

OpenStudy (anonymous):

less than or equal too. right before 25. Apologies

OpenStudy (aripotta):

and what about the choices?

OpenStudy (aripotta):

\[7-9x-(x+12)\le25\]

OpenStudy (anonymous):

-3x > 24 -3x < 4 -3x > -4 8 < x

OpenStudy (aripotta):

ok well first, you wanna distribute the - into the parentheses

OpenStudy (aripotta):

the - before the parentheses is understood to be -1 so you're really just changing the signs for everything in the parentheses

OpenStudy (aripotta):

\[7 - 9x -x - 12 \le 25\] then combine like terms

OpenStudy (anonymous):

Ok. Im fallowing

OpenStudy (aripotta):

so -9x and -x are like terms, as well as 7 and -12. combining them gives us \[-10x - 5 \le 25 \]

OpenStudy (anonymous):

so then we would go −10x−5+5≤ 25 +5?

OpenStudy (aripotta):

so now we start the process of getting x by itself. to do that we're going to the opposite operations. -5 is being subtracted, so we're going to add 5 to both sides. the 5 will be cancelled on the left side and make the 25 on the right, a 30

OpenStudy (anonymous):

−10x−5+5≤25+5?

OpenStudy (anonymous):

so then it would be: −10x≤20?

OpenStudy (aripotta):

it would be \[-10x \le 30\]

OpenStudy (anonymous):

Oh right! duh

OpenStudy (aripotta):

and since -10 is being multiplied to x, and the opposite of multiplication is division, we're going to divide both sides by -10

OpenStudy (anonymous):

then we devide both sides by -10?

OpenStudy (aripotta):

and remember, when you multiply or divide an inequality by a negative number, you have to flip the inequality sign

OpenStudy (anonymous):

x ≤ -3?

OpenStudy (aripotta):

so \[x \ge -3\]

OpenStudy (anonymous):

so how does that form the answer?

OpenStudy (anonymous):

theres still the -4

OpenStudy (aripotta):

what -4??

OpenStudy (anonymous):

Sorry looking at the wrong problem :P. Thank you soo much!

OpenStudy (aripotta):

oh lol

OpenStudy (anonymous):

cheers

OpenStudy (aripotta):

no problemo :)

OpenStudy (aripotta):

good luck on your other questions